obtain the expression for Fringe width of a bright fringe obtained in a interference experiment of Young.

Grade:12

1 Answers

hemema kurimidde iiit basar
6 years ago
if in youngs SINGLE slits experiment let the gap betwwen slits is ‘a’ then the distance between screen and slit is say R and draw a line exactly to the center of bright fring and consider a bright fring which is at a distance of ym and draw a line from the slit so we get a triangle 
as we know brigh fring is due to constructive interference so from path diference we get
Dsin(∂)=lamda×m
sin∂=lamda×m÷a
from our triangle ;tan∂=yR
AS ∂ IS VERY SMALL sin∂~tan∂
lamda×m÷d=ym÷R
ym=lamda×m×R÷a
assume for a bright at a distance of ym+1
we get ym+1=(m+1)lamda×R÷d
by subtracting ym+1-ym=lamda×R÷a